# SIDA: Social Media Image Deepfake Detection, Localization and Explanation with Large Multimodal Model

### v2

@@ 5 Experiments
| AntifakePrompt [5] | 2024 | 64.8 | 78.6 | 93.8 | 96.8 | 30.8 | 47.2 | 63.1 | 69.3 |
| CnnSpott [17] | 2021 | 79.8 | 88.7 | 39.5 | 56.6 | 6.9 | 12.9 | 42.1 | 69.6 |
| FreDect [18] | 2020 | 83.7 | 91.1 | 16.8 | 28.8 | 11.9 | 21.3 | 37.4 | 23.4 |
| Fusing [27] | 2022 | 85.1 | 92.0 | 34.0 | 50.7 | 2.7 | 5.3 | 40.1 | 29.1 |
| Gram-Net [39] | 2020 | 70.1 | 82.4 | 93.5 | 96.6 | 0.8 | 1.6 | 55.0 | 58.0 |
| UnivFD [48] | 2023 | 68.0 | 67.4 | 62.1 | 87.5 | 64.0 | 85.3 | 64.0 | 85.3 |
| LGrad [61] | 2023 | 64.8 | 78.6 | 83.5 | 91.0 | 6.8 | 12.7 | 51.8 | 55.5 |
| LNP [2] | 2023 | 71.2 | 83.2 | 91.8 | 95.7 | 2.9 | 5.7 | 55.2 | 58.2 |
| SIDA-7B | 2024 | 89.1 | 91.0 | 98.7 | 98.6 | 91.2 | 91.0 | 93.5 | 93.5 |
@@ 5.4 Test on Other Benchmark
| CNNSpot [17] | 87.8 | 88.4 | 28.4 | 44.2 | 40.6 | 43.3 |
| Gram-Net [39] | 62.8 | 54.1 | 78.8 | 88.1 | 67.4 | 79.4 |
| Fusing [27] | 87.7 | 86.1 | 15.5 | 27.2 | 40.4 | 36.5 |
| LNP [2] | 63.1 | 67.4 | 56.9 | 72.5 | 58.2 | 68.3 |
| UnivFD [48] | 89.4 | 88.3 | 44.9 | 61.2 | 53.9 | 60.7 |
| AntifakePrompt [5] | 91.3 | 92.5 | 89.3 | 91.2 | 90.6 | 91.2 |
| SIDA-7B | 92.9 | 93.1 | 90.7 | 91.0 | 91.8 | 92.4 |

### v3

@@ 5 Experiments
| AntifakePrompt [5] | 2024 | 64.8 | 78.6 | 93.8 | 96.8 | 30.8 | 47.2 | 63.1 | 74.2 |
| CnnSpott [17] | 2021 | 79.8 | 88.7 | 39.5 | 56.6 | 6.9 | 12.9 | 42.1 | 52.7 |
| FreDect [18] | 2020 | 83.7 | 91.1 | 16.8 | 28.8 | 11.9 | 21.3 | 37.4 | 47.0 |
| Fusing [27] | 2022 | 85.1 | 92.0 | 34.0 | 50.7 | 2.7 | 5.3 | 40.6 | 49.3 |
| Gram-Net [39] | 2020 | 70.1 | 82.4 | 93.5 | 96.6 | 0.8 | 1.6 | 54.8 | 60.2 |
| UnivFD [48] | 2023 | 68.0 | 67.4 | 62.1 | 87.5 | 64.0 | 85.3 | 64.7 | 80.0 |
| LGrad [61] | 2023 | 64.8 | 78.6 | 83.5 | 91.0 | 6.8 | 12.7 | 51.7 | 60.7 |
| LNP [2] | 2023 | 71.2 | 83.2 | 91.8 | 95.7 | 2.9 | 5.7 | 55.3 | 61.5 |
| SIDA-7B | 2024 | 89.1 | 91.0 | 98.7 | 98.6 | 92.7 | 91.0 | 93.5 | 93.5 |
@@ 5.4 Test on Other Benchmark
| CNNSpot [17] | 87.8 | 88.4 | 28.4 | 44.2 | 58.1 | 66.3 |
| Gram-Net [39] | 62.8 | 54.1 | 78.8 | 88.1 | 70.8 | 71.1 |
| Fusing [27] | 87.7 | 86.1 | 15.5 | 27.2 | 51.6 | 56.6 |
| LNP [2] | 63.1 | 67.4 | 56.9 | 72.5 | 60.0 | 70.0 |
| UnivFD [48] | 89.4 | 88.3 | 44.9 | 61.2 | 67.2 | 74.8 |
| AntifakePrompt [5] | 91.3 | 92.5 | 89.3 | 91.2 | 90.3 | 91.8 |
| SIDA-7B | 92.9 | 93.1 | 90.7 | 91.0 | 91.8 | 92.0 |